Question9 A proton in a particular electric field experiences an electric force whose magnitude is equal to the weight of the proton. Determine the magnitude of the electric field at the location of the proton hethe magitude of the electric Select the correct answer O 10- N/C O 10-3N/c O 10-4N/c O 10 N/C O 10-6N/c
Answer #1

Electric force in electric field is given by:

Fe = q*E

q = charge on proton = +e = 1.6*10^-19 C

Now given that

Fe = Weight of proton

Fe = m*g

q*E = m*g

E = m*g/q

m = mass of proton = 1.67*10^-27 kg

g = 9.8 m/sec^2

So Using these values:

E = 1.67*10^-27*9.8/(1.6*10^-19)

E = 1.0*10^7 N/C

E = 10^-7 N/C

Correct option is D.

Please Upvote.